# Infrared-Assisted Single-Stage Framework for Joint Restoration and Fusion of Visible and Infrared Images under Hazy Conditions

### v2

@@ I Introduction
In recent years, significant progress has been made in IR-VIS image fusion[1, 2, 3, 4, 5]; however, existing methods generally assume that the input visible images are of good visual quality. In hazy conditions, visible images are affected by haze, resulting in unclear imagery, which makes it difficult for these methods to generate clear, haze-free fusion results. Traditional approaches typically address this issue by first applying a dehazing algorithm to the hazy image and then fusing the dehazed image with the infrared image, as shown in Fig. 1(a). Although this two-stage strategy is feasible, it fails to integrate the dehazing and fusion tasks into a unified framework for joint training, making it challenging to balance the relationship between the two tasks. While dehazed images may show good dehazing performance, they are not always optimal for subsequent fusion tasks. Additionally, the two-stage process of dehazing followed by fusion involves different methodologies, reducing the model’s compactness.
@@
  • Innovative Joint Framework. An infrared-assisted joint learning framework for IR-VIS image fusion in hazy conditions is proposed, enabling collaborative training of dehazing and fusion tasks in a single stage. Compared to traditional two-stage methods, this framework more efficiently leverages the correlation between dehazing and fusion tasks, allowing them to complement each other and achieve a balanced outcome.
@@
  • Prompt Generation Module and Multi-Stage Fusion. A prompt generation module is designed to create a prompt selection matrix leveraging non-shared information to intelligently select and embed prompts that assist visible image restoration under hazy conditions. This approach not only enhances the restoration of visible image features but also minimizes the adverse impact of modality differences between IR-VIS features. Furthermore, a multi-stage feature fusion mechanism based on prompt embeddings is proposed to facilitate feature compensation and refinement, further improving the fusion result.
  • Efficient and Compact Model Structure. The proposed method adopts a single-stage processing framework, making it more compact than the traditional two-stage dehazing and fusion approach, thereby facilitating a better balance between the two tasks. Furthermore, the method efficiently coordinates the dehazing and fusion, enabling the restored image features to more effectively support the generation of clear, haze-free, high-quality fusion results. This design not only enhances the fusion effect but also improves the model’s applicability.
@@
II Related Work
II-A Typical Fusion Methods
In IR-VIS image fusion, traditional methods based on multi-scale transforms and sparse representations [9, 10, 11, 12, 13] remain relevant. However, deep learning-based techniques have become mainstream. These methods can be categorized into three types: Convolutional Neural Network (CNN)-based methods [14, 15, 16, 17], hybrid CNN-Transformer methods [18, 19, 20, 21], and Generative Adversarial Network (GAN)-based methods [22, 23, 24, 25, 26, 27, 4]. CNN-based methods extract features from input images and perform fusion using specialized modules, enhancing image details and contrast. However, CNNs are limited in modeling long-range dependencies, which impacts fusion quality in complex scenes. In contrast, Transformers excel at capturing long-range dependencies but struggle with local details and edges. Hybrid methods, such as AFT [19], YDTR [20], and HitFusion [21], combine CNN with Transformer to model local and global information, improving fusion performance.
II-B Methods Under Complex Imaging Conditions
To address the degradation of visible images under complex conditions, a decomposition-based, interference-aware fusion method was proposed [34], which effectively fuses degraded IR-VIS images, including noisy, overexposed, and snowy images, but does not address hazy conditions. To address this, AWFusion [6] introduces a clear feature prediction module based on the atmospheric scattering model, enabling dehazing capabilities. However, AWFusion also considers other weather conditions like snow and rain, which reduces its effectiveness in hazy conditions. To balance these tasks, Text-IF [7] uses text guidance for fusion by generating modulation parameters to control cross-attention outputs, but it is not specifically designed for hazy IR-VIS fusion, resulting in suboptimal performance. Additionally, the need for text input limits its convenience. VIFNet [35] restores hazy images using infrared images, but it focuses only on dehazing and does not address fusion. In contrast, this paper specifically focuses on IR-VIS fusion in hazy conditions to achieve clear, haze-free results.
III The Proposed Method
III-A Overview
III-B Prompt Generation
@@
As shown in Fig. 2, in the PGM, we first utilize an encoder constructed with Restormer [36] to perform feature encoding on the input registered IR-VIS images $\{{\bm{I}}_{ir},{\bm{I}}_{vi}\}$. As depicted in Fig. 2(a), Restormer consists of a self-attention layer and a feed-forward network layer. The features output by the Restormer encoder are denoted as ${{\bm{F}}_{vi}}\in\mathbb{R}^{C\times H\times W}$ and ${{\bm{F}}_{ir}}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ represent the number of channels, height, and width of the features, respectively. Additionally, in this module, the features of the hazy visible image ${\bm{I}}_{vi}$ and the infrared image ${\bm{I}}_{ir}$ are processed by:
@@
to remove the shared information and highlight the unique information. The resulting difference $\bm{F}_{vi-ir}$ is then fed into a weight prediction network composed of Convolutional (Conv) layer, GAP, Linear layer, and Softmax, resulting in a weight matrix ${{\bm{W}}_{p}}\in\mathbb{R}^{C\times H\times W}$ for selecting prompt information from the prompt pool $\bm{P}_{ir}$. At this stage, the prompt embedding generated for compensating the infrared image features can be represented as:
${{\bm{\hat{P}}}_{ir}}={\rm{Conv}}_{3\times 3}\left({\rm{Conv}}_{1\times 1}\left({{\bm{W}}_{p}}\odot{{\bm{P}}_{ir}}\right)\right)$ (2)
@@
where ${\bm{P}_{ir}}\in\mathbb{R}^{C\times H\times W}$ is a learnable prompt pool, ${\rm{Conv}}_{3\times 3}$ and ${\rm{Conv}}_{1\times 1}$ denote $3\times 3$ and $1\times 1$ Conv layers, respectively. The resulting ${\bm{\hat{P}}_{ir}}$ is then fed, along with the infrared image features ${\bm{F}_{ir}}$, into the Prompt Embedding Block (PEB) to obtain the features ${\bm{\hat{F}}_{ir}}$ for compensating the dense haze regions in the visible image ${\bm{I}}_{vi}$.
III-C Feature Restoration Assisted by Infrared Image
@@
To effectively utilize the information provided by the infrared image $\bm{I}_{ir}$ for restoring features in haze regions, we design the IA-FRM. As shown in Fig. 2, when restoring features in hazy images, regions with higher haze density should receive more focus. Therefore, it is essential to estimate the haze density in the input images. To achieve this, we adopt the method from [8] to estimate the haze density in the input image features, as illustrated in Fig. 2(c). Specifically, the dark channel information of ${{\bm{F}}_{vi}}$ is obtained through the Dark Channel Generation (DCG) operation. The resulting dark channel information, along with the feature ${\tilde{\bm{F}}}_{vi}$, is then fed into the Atmospheric Light Estimation (ALE) module to predict the atmospheric light, denoted as $\bm{A}$. Here, ${\tilde{\bm{F}}}_{vi}$ is obtained by applying Channel Average Pooling (CAP) to ${{\bm{F}}_{vi}}$, followed by channel replication. Based on $\bm{A}$ and ${\tilde{\bm{F}}}_{vi}$, we can obtain the initial transmission map for each pixel location through the Get Haze Density (GHD) block, as shown in Eq. (3):
@@
${\bm{T}}=\mathbf{1}-\omega\cdot\text{DCG}\left(\tilde{\bm{F}}_{vi}\odot\bm{A}^{-1}\right)$ (3)
@@
where $\omega$ is a constant that adjusts the effect of the DCG prediction, set to 0.95 as suggested in [8]; $\mathbf{1}\in\mathbb{R}^{1\times H\times W}$ is a matrix of ones; and ${{\bm{A}}^{-1}}$ represents the element-wise reciprocal of the atmospheric light $\bm{A}$, which is broadcast along the channel dimension to match the dimension of ${{\bm{\tilde{F}}}_{vi}}$.
@@
Then, in the GHD block, the transmission map $\bm{T}$ is refined using a guided filter, yielding the refined transmission map $\bm{T}^{\prime}$. Since the transmission map is inversely proportional to the haze density, we estimate the haze density $\bm{H}$ through the following equation:
${\bm{H}}=\mathbf{1}-{\bm{T}}^{\prime}$ (4)
@@
In $\bm{H}$, the higher the haze density, the larger the corresponding value. Therefore, $\bm{H}$ is used to select the information for restoring $\bm{F}_{vi}$ from $\bm{\hat{F}}_{ir}$. In this process, we use $\left({\mathbf{1}-{\bm{H}}}\right)$ to remove the information from regions with heavier haze in ${\bm{F}}_{vi}$ and replace it with corresponding information from $\bm{\hat{F}}_{ir}$ to assist in restoring these regions. The specific process can be formulated as:
${\bm{\hat{F}}}_{vi}=TF\left({\bm{\hat{F}}}_{ir}\odot{\bm{H}}+{\bm{F}}_{vi}\odot\left(\mathbf{1}-{\bm{H}}\right)+{\bm{F}}_{vi}\right)$ (5)
@@
where $TF$ denotes Transformer block. To ensure the quality of ${\bm{\hat{F}}}_{vi}$, it is passed through a $3\times 3$ Conv to obtain the dehazed image ${\bm{\hat{I}}}_{vi}$, and $L_{1}$-loss is used to optimize the network:
$\ell_{1}=\left\|{\bm{\hat{I}}}_{vi}-{\bm{I}}_{vi,gt}\right\|_{1}$ (6)
@@
Here, ${\bm{I}}_{vi,gt}$ represents the corresponding ground truth haze-free visible image.
III-D Multi-stage Prompt Embedding Fusion
@@
IV Experiments
IV-A Experimental Configurations
@@
Metrics. To objectively evaluate the fusion performance of different methods, we adopt five commonly used image quality assessment metrics: Mutual Information ($Q_{MI}$) [39], Gradient-based Fusion Performance ($Q_{AB/F}$) [40], Chen-Varshney Metric ($Q_{CV}$) [41], Sum of Correlation of Differences ($Q_{SCD}$) [42], and Visual Information Fidelity ($Q_{VIF}$) [43]. These metrics are used to assess the quality of the fusion results, with clear source images (without haze) as reference images when necessary. Additionally, to evaluate the perceptual quality of the dehazing effects within the fusion results, we employ the Perceptual Index ($Q_{PI}$) [44], Natural Image Quality Evaluator ($Q_{NIQE}$) [45], and Spatial Frequency ($Q_{SF}$) [46]. According to the evaluation criteria, lower $Q_{CV}$, $Q_{PI}$, and $Q_{NIQE}$ values indicate better fusion performance, while higher values for the remaining metrics signify improved quality.
IV-B Implementation Details
All experiments are conducted using the PyTorch framework on a single 24GB NVIDIA GeForce RTX 4090 GPU. During training, images are randomly cropped to $256\times 256$ patches, with data augmentation techniques such as horizontal and vertical flipping applied. The model is trained for a total of 300 epochs, using a batch size of 6 and the AdamW optimizer [47]. The initial learning rate is set to $2\times 10^{-4}$ and is gradually reduced to $2\times 10^{-6}$ following a cosine annealing schedule.
IV-C Comparison with State-of-the-art Methods
In order to verify the effectiveness of our method, we compare it with two existing methods. The first methodology involves initially applying advanced image dehazing algorithms to remove haze from the visible images, followed by fusing the dehazed images with the infrared images. For this purpose, we select the latest and most effective dehazing methods, namely DIACMP [48] and Dehazeformer [49]. Next, we apply representative IR-VIS image fusion methods, such as MLFusion [33], U2Fusion [38], LRRNet [16], ALFusion [18], TIMFusion [50], and MRFS [51], to fuse the dehazed visible images with the infrared images. The second methodology employs the Text-IF method [7], which directly restores and fuses hazy images with the assistance of text information.

### v3

@@ 1 Introduction
In recent years, the emergence of deep learning has rapidly advanced numerous areas within computer vision[10, 39, 54, 8], and infrared–visible (IR–VIS) image fusion has achieved significant progress[7, 51]; however, existing methods generally assume that the input visible images are of good visual quality. In hazy conditions, visible images are affected by haze, resulting in unclear imagery, which makes it difficult for these methods to generate clear, haze-free fusion results. Traditional approaches typically address this issue by first applying a dehazing algorithm to the hazy image and then fusing the dehazed image with the infrared image, as shown in Fig. 1(a). Although this two-stage strategy is feasible, it fails to integrate the dehazing and fusion tasks into a unified framework for joint training, making it challenging to balance the relationship between the two tasks. While dehazed images may show good dehazing performance, they are not always optimal for subsequent fusion tasks. Additionally, the two-stage process of dehazing followed by fusion involves different methodologies, reducing the model’s compactness.
@@
  • 1.
    We propose an infrared-assisted single-stage framework for IR–VIS fusion under hazy conditions, jointly optimizing dehazing and fusion. Compared with two-stage pipelines, it better exploits their correlation to achieve more balanced results. The compact design efficiently coordinates both processes, enabling restored features to produce clearer, haze-free, high-quality fused images and improving practical applicability.
  • 2.
@@
    We design a prompt generation module that leverages non-shared information to construct a prompt selection matrix, enabling adaptive prompt selection and embedding to assist visible image restoration in hazy scenarios and alleviate modality discrepancies between infrared and visible features. In addition, we propose a multi-layer fusion mechanism based on prompt embeddings for feature compensation and progressive refinement, further improving the fusion quality.
  • 3.
    Extensive experiments conducted on the MSRS, M3FD, and RoadScene datasets demonstrate that the proposed framework, while maintaining relatively low model complexity, achieves superior or comparable performance to a wide range of recent fusion and dehazing methods in terms of both subjective visual quality and objective evaluation metrics.
2 Related Work
2.1 Typical Fusion Methods
In IR-VIS image fusion, traditional methods based on multi-scale transforms and sparse representations [18, 24, 52, 16] remain relevant. However, deep learning-based techniques have become mainstream. These methods can be categorized into three types: Convolutional Neural Network (CNN)-based methods [17, 25, 19], hybrid CNN-Transformer methods [15, 3, 34, 5], and Generative Adversarial Network (GAN)-based methods [28, 37, 27, 23, 20]. CNN-based methods extract features from input images and perform fusion using specialized modules, enhancing image details and contrast. However, CNNs are limited in modeling long-range dependencies, which impacts fusion quality in complex scenes. In contrast, Transformers excel at capturing long-range dependencies but struggle with local details and edges. Hybrid methods, such as AFT [3], YDTR [34], and HitFusion [5], combine CNN with Transformer to model local and global information, improving fusion performance.
2.2 Methods Under Complex Imaging Conditions
To address the degradation of visible images under complex conditions, a decomposition-based and interference-aware fusion method was proposed in [21], which is capable of handling multiple degradations such as noise, overexposure, and snow, but does not involve hazy scenarios. To tackle haze, AWFusion [22] introduces a clear feature prediction module based on the atmospheric scattering model, thereby enabling dehazing capability. However, AWFusion simultaneously considers various weather conditions such as snow and rain, which reduces its effectiveness specifically in hazy scenarios. To balance multiple tasks, Text-IF [42] employs text guidance and generates modulation parameters to control cross-modal attention outputs. Nevertheless, Text-IF is not specifically designed for hazy IR–VIS fusion, thus showing limited performance in such conditions, and the requirement of textual input also limits its practicality. OmniFuse [47] and Text-DiFuse [46] explicitly couple diffusion models with multimodal fusion, removing compound degradations and integrating information either in the latent space or during the diffusion process, while achieving controllable enhancement of specific semantic targets via text modulation. However, these approaches mainly focus on general compound degradations and lack targeted modeling for hazy IR–VIS fusion, while their dependence on textual or detection modules introduces additional cost for real deployment. Deno-IF [36] addresses multi-noise infrared–visible fusion and obtains high-quality results from noisy inputs through unsupervised denoising and feature restoration, but mainly focuses on noise degradation rather than hazy conditions. In addition, CFMW [35] achieves joint optimization of multi-weather removal and visible–infrared object detection through a weather-removal diffusion model and a cross-modal fusion Mamba architecture. However, CFMW mainly targets the detection task rather than generating high-quality hazy fusion images. VIFNet [43] restores hazy images using infrared guidance, but focuses only on dehazing and does not perform fusion. In contrast, this paper specifically targets IR–VIS fusion under hazy conditions and aims to obtain clear, haze-free fusion results.
3 Proposed Method
3.1 Overview
@@
3.2 Prompt Generation
@@
As shown in Fig. 2, in the PGM, we first utilize an encoder constructed with Restormer [45] to perform feature encoding on the input registered IR-VIS images $\{\bm{I}_{ir},\bm{I}_{vi}\}$. As depicted in Fig. 2(a), Restormer consists of a self-attention layer and a feed-forward network layer. The features output by the Restormer encoder are denoted as $\bm{F}_{vi}\in\mathbb{R}^{C\times H\times W}$ and $\bm{F}_{ir}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ represent the number of channels, height, and width of the features, respectively. Additionally, in this module, the features of the hazy visible image $\bm{I}_{vi}$ and the infrared image $\bm{I}_{ir}$ are processed by
@@
to remove the shared information and highlight the unique information. The resulting difference $\bm{F}_{vi-ir}$ is then fed into a weight prediction network composed of Convolutional (Conv) layer, GAP, Linear layer, and Softmax, resulting in a weight matrix $\bm{W}_{p}\in\mathbb{R}^{C\times H\times W}$ for selecting prompt information from the prompt pool $\bm{P}_{ir}$. We implement $\bm{P}_{ir}$ as a set of learnable prompts rather than a single fixed feature map:
$\bm{P}_{ir}=\big\{\mathcal{P}_{ir}^{l}\big\}_{l=1}^{L},$ (2)
Each $\mathcal{P}_{ir}^{l}$ is a learnable prompt. During training, the $L$ prompts are randomly initialized and jointly optimized together with the backbone network. Subsequently, these prompts are stacked into a learnable tensor, resulting in the prompt representation $\bm{P}_{ir}\in\mathbb{R}^{C\times H\times W}$.
At this stage, the prompt embedding generated for compensating the infrared image features can be represented as
$\hat{\bm{P}}_{ir}=\mathrm{Conv}_{3\times 3}\left(\mathrm{Conv}_{1\times 1}\left(\bm{W}_{p}\odot\bm{P}_{ir}\right)\right),$ (3)
@@
where $\mathrm{Conv}_{3\times 3}$ and $\mathrm{Conv}_{1\times 1}$ denote $3\times 3$ and $1\times 1$ Conv layers, respectively. The spatially varying weight matrix $\bm{W}_{p}$ predicted from $\bm{F}_{vi-ir}$ modulates the prompt tensor via element-wise multiplication, producing a content-adaptive prompt representation that is further refined by the subsequent convolutional layers to obtain $\hat{\bm{P}}_{ir}$. The resulting $\hat{\bm{P}}_{ir}$ is then fed, along with the infrared image features $\bm{F}_{ir}$, into the Prompt Embedding Block (PEB) to obtain the features $\hat{\bm{F}}_{ir}$ for compensating the dense haze regions in the visible image $\bm{I}_{vi}$.
3.3 Feature Restoration Assisted by Infrared Image
@@
To effectively utilize the information provided by the infrared image $\bm{I}_{ir}$ for restoring features in haze regions, we design the IA-FRM. As shown in Fig. 2, when restoring features in hazy images, regions with higher haze density should receive more focus. Therefore, it is essential to estimate the haze density in the input images. To achieve this, we adopt the method from [12] to estimate the haze density in the visible branch, as illustrated in Fig. 2(c).
Let $\bm{F}_{vi}\in\mathbb{R}^{C\times H\times W}$ denote the visible feature map output by the encoder. In the HDE, we first apply Channel Average Pooling (CAP) to $\bm{F}_{vi}$ to obtain a single-channel feature map
$\bar{\bm{F}}_{vi}=\mathrm{CAP}(\bm{F}_{vi})\in\mathbb{R}^{1\times H\times W},$ (4)
and then replicate it along the channel dimension (the “Copy” operation in Fig. 2(c)) to form a pseudo-RGB feature $\tilde{\bm{F}}_{vi}\in\mathbb{R}^{C\times H\times W}$. Based on $\tilde{\bm{F}}_{vi}$, the Dark Channel Generation (DCG) module computes the dark channel
$D(i,j)=\min_{c\in\{1,2,3\}}\left(\min_{(u,v)\in\Omega(i,j)}\bm{\tilde{F}}_{vi}^{c}(u,v)\right),$ (5)
where $\Omega(i,j)$ denotes a local window centered at pixel $(i,j)$ and $\bm{\tilde{F}}_{vi}^{c}$ is the $c$-th channel of $\tilde{\bm{F}}_{vi}$. The dark channel $D$ highlights regions that are heavily affected by haze.
Figure 3: (a)Illustration of the Atmospheric Light Estimation (ALE) module.(b)Illustration of the Get Haze Density (GHD) module.
In the Atmospheric Light Estimation (ALE) module, As shown in Fig. 3(a),we first select the top $\rho\%$ brightest pixels in $D$ (set to $\rho=0.1$ as suggested in [12]) and denote their index set as $\Omega_{A}$. The global atmospheric light $\bm{A}\in\mathbb{R}^{3}$ is then estimated by averaging the corresponding intensities of $\tilde{\bm{F}}_{vi}$:
$\bm{A}=\frac{1}{|\Omega_{A}|}\sum_{(i,j)\in\Omega_{A}}\tilde{\bm{F}}_{vi}(i,j).$ (6)
In the Get Haze Density (GHD) block, As shown in Fig. 3(b), the estimated atmospheric light $\bm{A}$ is first broadcast along the channel dimension and used to normalize the visible features. By combining $\bm{A}$ with $\tilde{\bm{F}}_{vi}$, we obtain the initial transmission map
@@
$\bm{T}=\mathbf{1}-\omega\cdot\mathrm{DCG}\!\left(\tilde{\bm{F}}_{vi}\odot\bm{A}^{-1}\right),$ (7)
@@
where $\omega$ is a constant that adjusts the effect of the DCG prediction (set to $0.95$ as suggested in [12]); $\mathbf{1}\in\mathbb{R}^{1\times H\times W}$ is a matrix of ones; $\bm{A}^{-1}$ represents the element-wise reciprocal of $\bm{A}$; and $\odot$ denotes the Hadamard product.
@@
Next, the coarse transmission map $\bm{T}$ is refined using a guided filter, yielding the refined transmission map $\bm{T}^{\prime}$. Since the transmission map is inversely proportional to the haze degree, the haze density map $\bm{H}$ is estimated as
$\bm{H}=\mathbf{1}-\bm{T}^{\prime}.$ (8)
@@
Pixels with higher values in $\bm{H}$ correspond to regions with heavier haze. Therefore, $\bm{H}$ is used to select the information for restoring $\bm{F}_{vi}$ from $\hat{\bm{F}}_{ir}$. In this process, we use $\left(\mathbf{1}-\bm{H}\right)$ to suppress the severely degraded regions in $\bm{F}_{vi}$ and replace them with the corresponding information from $\hat{\bm{F}}_{ir}$. The specific process can be formulated as
$\hat{\bm{F}}_{vi}=TF\!\left(\hat{\bm{F}}_{ir}\odot\bm{H}+\bm{F}_{vi}\odot\left(\mathbf{1}-\bm{H}\right)+\bm{F}_{vi}\right),$ (9)
@@
where $TF$ denotes the Transformer block. To ensure the quality of $\hat{\bm{F}}_{vi}$, it is passed through a $3\times 3$ convolution to obtain the dehazed image $\hat{\bm{I}}_{vi}$, and an $L_{1}$ loss is used to optimize the network:
$\ell_{1}=\left\|\hat{\bm{I}}_{vi}-\bm{I}_{vi,gt}\right\|_{1},$ (10)
@@
where $\bm{I}_{vi,gt}$ represents the corresponding ground-truth haze-free visible image.
3.4 Multi-stage Prompt Embedding Fusion
@@
4 Experiments
4.1 Experimental Configurations
@@
Metrics. To objectively evaluate the fusion performance of different methods, we adopt five commonly used image quality assessment metrics: Mutual Information ($Q_{MI}$) [30], Gradient-based Fusion Performance ($Q_{AB/F}$) [40], Chen-Varshney Metric ($Q_{CV}$) [4], Sum of Correlation of Differences ($Q_{SCD}$) [1], and Visual Information Fidelity ($Q_{VIF}$) [11]. These metrics are used to assess the quality of the fusion results, with clear source images (without haze) as reference images when necessary. Additionally, to evaluate the perceptual quality of the dehazing effects within the fusion results, we employ the Perceptual Index ($Q_{PI}$) [2], Natural Image Quality Evaluator ($Q_{NIQE}$) [29], and Spatial Frequency ($Q_{SF}$) [9]. Specifically, $Q_{PI}$ and $Q_{NIQE}$ are widely used no-reference naturalness-based perceptual metrics that reflect the overall visual quality and naturalness of restored images, while $Q_{SF}$ measures spatial frequency and thus reflects edge sharpness and the richness of spatial details. According to the evaluation criteria, lower $Q_{CV}$, $Q_{PI}$, and $Q_{NIQE}$ values indicate better fusion performance, while higher values for the remaining metrics signify improved quality.
4.2 Implementation Details
All experiments are conducted using the PyTorch framework on a single 24GB NVIDIA GeForce RTX 4090 GPU. In our implementation, the entire network is trained in an end-to-end manner under the joint restoration and fusion loss, so that both branches can be optimized collaboratively. During training, images are randomly cropped to $256\times 256$ patches, with data augmentation techniques such as horizontal and vertical flipping applied. The model is trained for a total of 300 epochs, using a batch size of 6 and the AdamW optimizer [13]. The initial learning rate is set to $2\times 10^{-4}$ and is gradually reduced to $2\times 10^{-6}$ following a cosine annealing schedule.
4.3 Comparison with State-of-the-art Methods
In order to verify the effectiveness of our method, we compare it with two existing methods. The first methodology involves initially applying advanced image dehazing algorithms to remove haze from the visible images, followed by fusing the dehazed images with the infrared images. For this purpose, we select the latest and most effective dehazing methods, namely DIACMP [49] and Dehazeformer [31]. Next, we apply representative IR-VIS image fusion methods, such as MLFusion [14], U2Fusion [38], LRRNet [19], ALFusion [15], TIMFusion [26], MRFS [48], SHIP [53] and FreeFusion [50], to fuse the dehazed visible images with the infrared images. The second methodology employs the Text-IF method [42], which directly restores and fuses hazy images with the assistance of text information.